OpenStudy (anonymous):

What is the name of this polygon? A. rectangle B. rhombus C. square D. trapezoid

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

umm i cant see your polygon

OpenStudy (anonymous):

does it helps

OpenStudy (anonymous):

Yes it is a trapezoid

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

A rectangle has an area of 36 square centimeters and a length of 9 centimeters. What is the width of the rectangle? A. 3 centimeters B. 27 centimeters C. 9 centimeters D. 4 centimeters

OpenStudy (anonymous):

can you help me

OpenStudy (anonymous):

okay thank you so much

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

Which correctly names the line in this figure? A.PQR B.Q C.K D.PR

OpenStudy (anonymous):

Which correctly names the line in this figure? A.PQR B.Q C.K D.PR

OpenStudy (anonymous):

ill think its d just bcuz PR u can say is the total or the complete line.

OpenStudy (anonymous):

*i

OpenStudy (anonymous):

so its D

OpenStudy (anonymous):

it makes sense PR it the whole line... im just saying

OpenStudy (anonymous):

*is

OpenStudy (anonymous):

okay that is fine thanks

OpenStudy (anonymous):

hope it helped

OpenStudy (anonymous):

Which pair of angles is a pair of alternate interior angles? VHS_PA_S1_07_L211_L311_UT_Q3.gif A. ∠3 and ∠8 B. ∠2 and ∠6 C. ∠4 and ∠6 D. ∠2 and ∠8

OpenStudy (anonymous):

There seems to be an "official"set of names for polygons, uniformly derived from the Greek (after 4 sides). After about 6 sides mathematicians usually refer to these polygons as n-gons. So your 13 sided polygon I would call a 13-gon.

OpenStudy (anonymous):

Which pair of angles is a pair of alternate interior angles? VHS_PA_S1_07_L211_L311_UT_Q3.gif A. ∠3 and ∠8 B. ∠2 and ∠6 C. ∠4 and ∠6 D. ∠2 and ∠8

OpenStudy (anonymous):

If r || s and m 5 = 65°, what is m 1? Geometry Basics, Part 1 A. 65° B. 115° C. 180° D. 25°

OpenStudy (anonymous):

A flag pole is 12 feet tall. It's shadow is 10 feet long. What is the approximate distance from the top of the flag pole to the tip of the pole's shadow? A. 22 feet B. 16 feet C. 11 feet D. 7 feet

OpenStudy (anonymous):

help

OpenStudy (anonymous):

C

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you are a smart girl

OpenStudy (anonymous):

Stephanie has been putting money in her savings account for the last year. The graph shows data about Stephanie's savings account over time. (coordinate graph) Which is a reasonable inference for the amount of money Stephanie had in her savings account at month 13? A. $600 B. $650 C. $750 D. $800

OpenStudy (anonymous):

The triangles shown are similar. What is the value of x? A. 13 B. 12 C. 15 D. 20

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

The triangles shown are similar. What is the value of x? A. 13 B. 12 C. 15 D. 20

OpenStudy (anonymous):

its over hear

OpenStudy (anonymous):

@

OpenStudy (anonymous):

its C; u have to times all the num. by 3

OpenStudy (anonymous):

okay that is it

OpenStudy (anonymous):

u see 4 x 3 is 12 , 3 x 3 is 9 5 x 3 is 15 hoped this helped

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

What is the surface area of the figure? A. 204 cm2 B. 300 cm2 C. 408 cm2 D. 360 cm2

OpenStudy (anonymous):

need help and fast

OpenStudy (anonymous):

A cone has a radius of 5 inches and a height of 18 inches. What is the exact volume of the cone? A. 150π in3 B. 300π in3 C. 450π in3 D. 540π in3

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

its A

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np. :-D

OpenStudy (anonymous):

Which is the equation of the line? A. 4x + y = 2 B. x – 4y = –8 C. x + 4y = –8 D. 4x – y = –2

OpenStudy (anonymous):

can someone help me

OpenStudy (anonymous):

What is the volume of the cylinder? Cylinder 18 feet high and a 7 foot radius A. 2268π ft3 B. 882π ft3 C. 126π ft3 D. 252π ft3

OpenStudy (anonymous):

help pleace

OpenStudy (anonymous):

The base area of a prism is 140 square feet. Its height is 6 feet. What is the volume of the prism? A. 146 ft3 B. 280 ft3 C. 420 ft3 D. 840 ft3

OpenStudy (anonymous):

The base area of a prism is 140 square feet. Its height is 6 feet. What is the volume of the prism? A. 146 ft3 B. 280 ft3 C. 420 ft3 D. 840 ft3

OpenStudy (nikato):

volume= base area x height

OpenStudy (anonymous):

so is 140*6?

OpenStudy (anonymous):

so i multiply it

OpenStudy (anonymous):

the answer that i got is 840 so is it D

OpenStudy (nikato):

correct

OpenStudy (anonymous):

thank you

OpenStudy (nikato):

no problem

OpenStudy (anonymous):

solve for x: x-3/2=-5 A. –16 B. –13 C. –7 D. –4

OpenStudy (anonymous):

help!!!

OpenStudy (nikato):

first multiply both sides by 2 then add 3 to both sides

OpenStudy (anonymous):

so like 3*3 and 2*2 or just 3*2

OpenStudy (anonymous):

how??

OpenStudy (nikato):